OpenStudy (anonymous):

Tony is 1/2 of a mile ahead of Laura on a jogging path. If Tony is running at 5 miles per hour and Laura is running at 7 miles per hour, how long will it take Laura to catch up with Tony? A. 1/2of an hour or 30 minutes B. 1/4 of an hour or 15 minutes C. 1/3of an hour or 20 minutes D. 1 hour or 60 minutes

OpenStudy (akshay_budhkar):

lol y dont u acknowledge by giving a medal?

OpenStudy (lilg132):

oh there we go i think hes got the message

OpenStudy (anonymous):

i am you idiot

OpenStudy (lilg132):

oops reported :p

OpenStudy (anonymous):

nerd

OpenStudy (akshay_budhkar):

lol more than medals learn to respect those who help.. u think i do it for medals? see we r here to help giving medals is a way of thanking us..

OpenStudy (lilg132):

call me what you want i know how to answer these questions where as your looking for answers instead of learning it for your self, just let me know how far you get in your exams. :)

OpenStudy (akshay_budhkar):

ok lets solve this together santi

OpenStudy (akshay_budhkar):

um you here santi?

OpenStudy (akshay_budhkar):

i feel he left :/

OpenStudy (lilg132):

dont think hes interested :p

OpenStudy (lilg132):

but anyways can you work it out for me im stuck on this :p

OpenStudy (akshay_budhkar):

ofcourse hehe see lets take the time as t hours so we have the distance travelled by person 1 is 5t and person 2 is 7t.. the difference in distance is 1/2.. so ur equation is something like 5t+1/2=7t

OpenStudy (akshay_budhkar):

so we have t=1/4 mile

OpenStudy (akshay_budhkar):

is that fine?

OpenStudy (lilg132):

5t + 1/2 = 7t 1/2 = 2t 1/4 = t ok got that

OpenStudy (akshay_budhkar):

:D

OpenStudy (lilg132):

ahh so the answer is 1/4 or 15 minutes :p i get it thanks